Solve.
Jeff takes 5 hours longer to build a fence than it takes Bill. When they work together, it takes them 6 hours. How long would it take Bill to do the job alone?

Answer by checkley77(12844)

THE FORMULA FOR THIS PROBLEM IS THE PRODUCT OF THEIR INDIVIDUAL HOURS DIVIDED BY THE SUM OF THEIR HOURS:
X(X+5)/(X+X+5)=6
X^2+5X/(2X+5)=6
X^2+5X=6(2X+5)
X^2+5X=12X+30
X^2+5X-12X-30=0
X^2-7X-30=0
(X-10)(X+3)=0
X-10=0
X=10 HOURS IS THE ANSWER FOR BILL'S TIME TO BUILD THE FENCE.
10+5=15 HOURS FOR JEFF TO BUILD THE FENCE.
PROOF
10*15/(10+15)=6
150/25=6
6=6